OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (3, –1) and is perpendicular to the line y = x + 1?

OpenStudy (anonymous):

Here solve for y y + 1 = -1(x-3)

OpenStudy (anonymous):

y=-1x-3 ?

OpenStudy (anonymous):

Not quite its y = -x +3 -1 can you simplify this?

OpenStudy (anonymous):

y=-x+2 ?

OpenStudy (anonymous):

Yes.. there you go!

OpenStudy (anonymous):

so thats it?

OpenStudy (anonymous):

wow that was easy thank you!

OpenStudy (anonymous):

yw =))